what am I doing wrong?
x+x+1+x+2+x+4+x+6=-2
4x+12=-2
4x=-14
x=-3.5
x+2=-1.5
x+4=0.5
x+6=2.5

Answer by swincher4391(1107)

Based on your initial setup, it seems you have miscounted.
x+x+1+x+2+x+4+x+6 = -2 would give you 5xs and (1+2+4+6) = 13
Hence,
5x + 13 = -2
5x = -15
+x+=+-3+